OpenStudy (anonymous):

simplify square root of 512k to 2nd power

OpenStudy (anonymous):

\[\sqrt{512,000^{2}}=512,000\]

OpenStudy (anonymous):

k being a variable

OpenStudy (mathmale):

I think of 512k as meaning 512000. Are you positive that k is a variable, and not shorthand for 1000?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k is variable

OpenStudy (anonymous):

sorry, if "k" is a variable, then 16k√2

OpenStudy (mathmale):

All right, then, k is to be considered a variable. Which do you mean: \[(\sqrt{512k})^{2},or \sqrt{(512k)^{2}}, or\] are these the same?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

√ an ^2 cancel out=>512k

OpenStudy (mathmale):

OK. Square 512k now. Once done, please find the square root of the result. Dalia's statement is valid; that method involves a lot less work.

OpenStudy (anonymous):

thank you